SOLVE:
5\8- 3\5 = x\10

Answer by stanbon(57356)

SOLVE:
5\8- 3\5 = x\10
----------------
Multiply thru by 40 to get:
25 - 24 = 4x
1 = 4x
x = 1/4
=============
Cheers,
Stan H.